A chmistry student needs 10L of a 12% acid solution. The laboratory has an ample amounts of 10% and 20% acid solutions. How many liters of each solution should the student mix?
Answer by longjonsilver(2297)

mixture questions are all about BEFORE and AFTER...they have to be equal.

Let x=amount of 10% acid to be added.
and so how much of the 20% acid? Leap of faith... it is everything else of the 10litres of the final amount surely? so (10-x)litres of the 20% acid.

--> 0.10*x + 0.20(10-x) is everything before mixing.

After mixing, we have 10 litres of a 12% acid --> 0.12*10

so 0.10*x + 0.20(10-x) = 0.12*10
0.1x + 2 - 0.2x = 1.2
-0.1x + 2 = 1.2
-0.1x = -0.8
x = -0.8/-0.1
x = 8 litres

and so we need 8 litres of the 10% acid and 2 litres of the 20% acid.

Easy? :-)

jon.